# The Behrens–Fisher problem revisited

### v1

@@ 2 Mellin–Barnes representation of $T$
Consider the Behrens–Fisher statistic $T=\frac{\bar{X}-\bar{Y}}{\sqrt{S_{1}^{2}/n_{1}+S_{2}^{2}/n_{2}}}$ with $\nu_{i}=n_{i}-1$ degrees of freedom, and recall the notation introduced in Section 1: $g=\frac{\sigma_{1}^{2}}{n_{1}}$, $h=\frac{\sigma_{2}^{2}}{n_{2}}$, $W_{1}=\frac{\nu_{1}S_{1}^{2}}{\sigma_{1}^{2}}$, $W_{2}=\frac{\nu_{2}S_{2}^{2}}{\sigma_{2}^{2}}$. Under $H_{0}:\mu_{1}=\mu_{2}$, we have $\bar{X}-\bar{Y}\left|\left(\sigma_{1}^{2},\sigma_{2}^{2}\right)\right.\sim\mathcal{N}\!\left(0,g+h\right)$, and this difference is independent of $(S_{1}^{2},S_{2}^{2})$. Hence, conditional on the random denominator, the statistic $T$ is merely a scaled normal ratio:
$T\left|(W_{1},W_{2})\right.\sim\frac{\mathcal{N}\left(0,g+h\right)}{\sqrt{\frac{gW_{1}}{\nu_{1}}+\frac{hW_{2}}{\nu_{2}}}}$ (11)
with the following pdf:
@@
where $W_{1}\sim\chi^{2}_{\nu_{1}}$, $W_{2}\sim\chi^{2}_{\nu_{2}}$, $W_{1}\perp W_{2}$, so their pdfs are
$f_{W_{i}}(w)$ $=$ $\frac{w^{\nu_{i}/2-1}\,e^{-w/2}}{2^{\nu_{i}/2}\Gamma(\nu_{i}/2)},\quad w>0,\quad i=1,2.$ (13)
There is no closed-form pdf for the weighted sum $\sqrt{gW_{1}/\nu_{1}+hW_{2}/\nu_{2}}$; this is the core obstacle of the Behrens–Fisher problem, so we keep the bivariate $(W_{1},W_{2})$ pair intact.
@@ 2.1 MB expansion of the square-root factor
$(x+y)^{\frac{1}{2}}$ $=$ $\frac{1}{2\pi i}\int\limits_{c-i\infty}^{c+i\infty}\frac{\Gamma(-s)\,\Gamma(\tfrac{1}{2}+s)}{\Gamma(\tfrac{1}{2})}x^{s}\,y^{1/2-s}\,ds,$ (16)
@@
The inseparable factor is now a weighted geometric mean $(g/\nu_{1})^{s}w_{1}^{s}(h/\nu_{2})^{1/2-s}w_{2}^{1/2-s}$ multiplied by a pures-s coefficient (the $\Gamma$-product). All $w_{1}$-dependence and $w_{2}$-dependence are split into separate powers. This makes the forthcoming integrations elementary.
Since absolutely convergent integrals may be rearranged, we move the MB contour outside:
@@
where $\alpha=\dfrac{t^{2}}{2(g+h)}$. With this development, (18) is precisely the single-contour formula. Two dimensions have become one; the price is the integral over a vertical line in the complex plane. From here, residue calculus yields an explicit series by closing the contour to the right: poles of $\Gamma(-s)$ at $s=0,1,\dots$ give the hypergeometric-type series, which terminates when either $\nu_{1}$ or $\nu_{2}$ is odd. Alternatively, before the MB step, one may collapse $(w_{1},w_{2})$ to $(S,U)$. But, the MB path is shorter: from (18), shift the contour left, pick the poles of $\Gamma(\tfrac{1}{2}+s)$, and recognize the resulting Beta function integral as Gauss’ hypergeometric, thereby facilitating a compact expression for $f_{T}(t)$ that becomes apparent in the sequel. More importantly, the integrand in (18) is exactly of the $\gamma$-product type that Ramanujan’s master theorem turns into asymptotic series without further work: expand the $\Gamma$-ratio around large $s$ to read off the $\lvert t\rvert\to\infty$ expansion coefficients directly, thereby yielding exact tail coefficients.
The MB approach works for all $g$, $h$. While Taylor series in $w_{1}/w_{2}$ converges only when one weight is smaller than the other, the MB integral is valid everywhere. It handles non–integer exponents naturally. In our case, the exponent is $\rho=\tfrac{1}{2}$; the MB representation is tailor–made for arbitrary complex $\rho$. Note that, MB aids interfacing with $\Gamma$-distributions: $\chi^{2}$ densities are special cases of $\Gamma$-densities. The MB splits $(x+y)^{\rho}$ into $\Gamma$-functions of $s$ times monomials in $x$ and yy—perfectly aligned with $\Gamma$ integrals. Furthermore, the MB integrals provide contour flexibility. We may choose to close left or right, whichever set of poles yields the nicer series (terminating, fastest convergence, etc).
@@ 2.2 Residue calculus on the MB contour
$\Gamma(-s)$ $s=0,1,2,\dots$ Right-hand ladder
$\Gamma(\tfrac{1}{2}+s)$ $s=-\tfrac{1}{2},-\tfrac{3}{2},-\tfrac{5}{2},\dots$ Left-hand ladder
$\Gamma\left(s+\tfrac{\nu_{1}}{2}\right)$ $s=-\tfrac{\nu_{1}}{2},-\tfrac{\nu_{1}}{2}-1,\dots$ Left ladder, terminates if $\nu_{1}$ even
$\Gamma\left(\tfrac{\nu_{2}}{2}-s+\tfrac{1}{2}\right)$ $s=\tfrac{\nu_{2}}{2}+\tfrac{1}{2}+k,\ k=0,1,\dots$ Right ladder, terminates if $\nu_{2}$ odd
@@
the factor $\Gamma\!\left(\tfrac{\nu_{2}}{2}-k+\tfrac{1}{2}\right)$ vanishes once $k>\tfrac{\nu_{2}}{2}-\tfrac{1}{2}$ if $\nu_{2}$ is an odd integer because the $\Gamma$-argument hits a non-positive integer. The same happens to $\Gamma\!\left(k+\tfrac{\nu_{1}}{2}\right)$ when $\nu_{1}$ is even and we close the contour to the left (alternative route). Thus the series terminates whenever either degrees of freedom parameter is odd—a property first noted by [2]. Collecting all right-hand residues and the front constants in (21) gives
@@
This series is valuable for the following reasons: The $\Gamma$-coefficients are rational in $\nu_{1},\nu_{2}$; algebra systems can integrate or differentiate $f_{T}(t)$ term-wise to obtain moments, mgf, cdf expansions, etc. Only a handful of terms are needed for double-precision accuracy when $t\lesssim 10$. Truncation after $m$ terms leaves a remainder $O(t^{-(\nu_{1}+\nu_{2}+1)+2m})$, giving controlled error bounds directly. Recognizing the ratio of $\Gamma$’s as Beta coefficients, (23) rearranges into a Gauss hypergeometric series $_{2}F_{1}$; summation leads to a closed form. If $g=h$, the factor $(g/h)^{k}$ becomes 1 and the $\Gamma$-triple collapses to a single $\Gamma$-ratio, giving the $t_{\nu_{1}+\nu_{2}}$ density via the duplication formula.
Summing (23) with the Euler-type identity for $_{2}F_{1}$ yields compact expression for $f_{T}(t)$ which we will provide in the next section. That closed form is preferred for large $\lvert t\rvert$, because the hypergeometric behaves smoothly, while the plain series may suffer cancellation. The same $\Gamma$-product that gives each residue also feeds Ramanujan’s master theorem. Essentially, we write
$\Psi(s)\,\Gamma(s+\nu_{1}/2)\,\Gamma(\nu_{2}/2-s+1/2)\equiv\Gamma(s)\,A(-s)$ (24)
and read off the coefficients $A(m)$ directly for the $|t|\to\infty$ expansion. When the degrees of freedom $\nu_{i}$, $i=1,2$, are both odd, we get a finite series that leads to an elementary rational function using duplication. When $\nu_{i}$ are both even, we get a hypergeometric series of type $_{2}F_{1}$ with integer parameters, reducible further to complete elliptic integrals (Carlson form) for certain balanced sample-size ratios. Reducing the contour integral into residues is central to our work that lets us transform an abstract complex integral into an explicit computational tool. It reveals hidden algebraic cancellations whenever a degree of freedom is integer, bridges to standard special-function framework (hypergeometric/Beta) and prepares the ground for both numerical evaluation and asymptotic analysis. Everything that follows—closed-form hypergeometric, tail coefficients, moment formulas—starts with the series (23). Without identifying and summing the $\Gamma(-s)$ residues, the Behrens–Fisher density would remain locked inside the contour integral; after residual analyses, it becomes fully transparent.
@@
The MB approach in the previous sections already gave a series. However, series are not always the best numerical object—convergence slows in the tails and severe cancellation can occur when $g\approx h$. A single real integral with Euler–Beta weight can sometimes be integrated analytically, producing a compact special-function formula (see [25, Section 2.3]). Most symbolic integrators (Mathematica, Maple, etc.) recognize Euler-type kernels and instantly reduce them to the classical Gauss function $_{2}F_{1}$. Hence, we reorganize the starting double integral so that all dependence on one variable disappears, leaving a 1-D integral of Euler–Beta form.
@@
where $\mathcal{C}$ is the set of all constants independent of $U$: $\mathcal{C}=\frac{\Gamma(c)(g/\nu_{1})^{-a}(h/\nu_{2})^{-b}}{2^{(a+b)}\Gamma(a)\Gamma(b)\sqrt{2\pi(g+h)}}$, with $a=\frac{\nu_{1}}{2}$, $b=\frac{\nu_{2}}{2}$, $c=\frac{\nu_{1}+\nu_{2}+1}{2}$. Upon closer inspection, we can write, with $\lambda=\frac{\nu_{1}h}{2\nu_{2}g}-\frac{\alpha h}{\nu_{2}}-1$,
$f_{T}(t)=\mathcal{C}(\nu_{2}/h)^{-c}\int_{0}^{1}\!u^{a-1}(1-u)^{b-1}(1+\lambda u)^{-c}du,$ (28)
By Euler’s integral representation of Gauss’ hypergeometric function (see [29, Chapter 15]), for $\Re c>0$, $\int_{0}^{1}\!u^{a-1}(1-u)^{b-1}(1+\lambda u)^{-c}du=B(a,b){}_{2}F_{1}(c,a;a+b;-\lambda)$, with $B(a,b)=\Gamma(a)\Gamma(b)/\Gamma(a+b)$ so that the Beta factor cancels the $\Gamma(a)\Gamma(b)$ already in $\mathcal{C}$. Letting $\rho(t)=\frac{g^{-1}-h^{-1}}{h^{-1}+t^{2}/(g+h)}$, and organizing the terms together yields
@@
where $\rho(t)$ is the $t$-adjusted version of $\lambda$ appears in the fully scaled hypergeometric formula.
@@
Natural starting point for Ramanujan’s theorem/saddle-point asymptotics Immediate access to Gauss–Kummer transformations for numerical conditioning.
@@
It is worth noting that the MB representation and Euler-type integral collapses represent two analytically distinct yet ultimately equivalent approaches to deriving the same result. For symbolic series manipulations, MB is the preferred approach. For compact numeric evaluation, Euler–hypergeometric is preferable. Table 2 summarizes the differences between these two approaches. Most scientific computing libraries (e.g., SciPy) support high-precision evaluation of $_{2}F_{1}$ with rigorous error control, enabling accurate computation of $f_{T}(t)$ for any parameters $(\nu_{1},\nu_{2},g,h,t)$. Interchanging the two samples, i.e., $(g,\nu_{1})\leftrightarrow(h,\nu_{2})$, leaves $f_{T}(t)$ invariant. This follows from the symmetry of $_{2}F_{1}(c,a;a+b;-\rho)$ in $(a,b)$ when $\rho\to-\rho$. In the equal-variance limit, as $g\to h$, we have $\rho\to 0$ and $_{2}F_{1}\to 1$, so (29) reduces to the Student $t$ density with $\nu_{1}+\nu_{2}$ degrees of freedom. In the large–sample limit, as $\nu_{1},\nu_{2}\to\infty$, the distribution converges to $\mathcal{N}(0,1)$. The asymptotic behavior of $_{2}F_{1}(c,a;a+b;-z)$ as $z\to\infty$ follows directly from Euler’s integral representation, yielding an $O\left(t^{-(\nu_{1}+\nu_{2}+1)}\right)$ decay. This matches the tail derived via complex–residue analysis. Term–by–term integration of (29) in $t$, under the substitution $t=\sqrt{g+h}\tan\theta$, converts the expression into a sum of Appell $F_{1}$ functions—making it amenable for numeric evaluation.
The one-dimensional collapse is the key that unlocks a single Gauss hypergeometric function for $f_{T}$. This form is: (i) analytically elegant—encapsulating all parameter dependence in five $\Gamma$’s and one $_{2}F_{1}$; (ii) numerically robust—ready-made for standard libraries and asymptotic expansions; (iii) conceptually clarifying—shows that the Behrens–Fisher density lives in the classical hypergeometric family, explaining why equal-variance data drop to Student-$t$ (the $_{2}F_{1}\!\equiv 1$ case) and why odd degrees of freedom collapse to elementary expressions (because the Gauss function terminates). Subsequent steps—cdf calculation, likelihood inference, saddle-point corrections, or benchmarking Welch–Satterthwaite—benefits from having this concise, one–line representation.

### v2

@@ 2 Mellin–Barnes representation of $T$
Consider the Behrens–Fisher statistic $T=\frac{\bar{X}-\bar{Y}}{\sqrt{S_{1}^{2}/n_{1}+S_{2}^{2}/n_{2}}}$ with $\nu_{i}=n_{i}-1$ degrees of freedom, and recall the notation introduced in Section 1: $g=\frac{\sigma_{1}^{2}}{n_{1}}$, $h=\frac{\sigma_{2}^{2}}{n_{2}}$, $W_{1}=\frac{\nu_{1}S_{1}^{2}}{\sigma_{1}^{2}}$, $W_{2}=\frac{\nu_{2}S_{2}^{2}}{\sigma_{2}^{2}}$. Under $H_{0}:\mu_{1}=\mu_{2}$, we have $\bar{X}-\bar{Y}\left|\left(\sigma_{1}^{2},\sigma_{2}^{2}\right)\right.\sim\mathcal{N}\!\left(0,g+h\right)$, and this difference is independent of $(S_{1}^{2},S_{2}^{2})$. Hence, conditional on the random denominator, the statistic $T$ is merely a scaled normal ratio $T\left|(W_{1},W_{2})\right.\sim\frac{\mathcal{N}\left(0,g+h\right)}{\sqrt{\frac{gW_{1}}{\nu_{1}}+\frac{hW_{2}}{\nu_{2}}}}$ with the following pdf:
@@
where $W_{1}\sim\chi^{2}_{\nu_{1}}$, $W_{2}\sim\chi^{2}_{\nu_{2}}$, $W_{1}\perp W_{2}$, so their pdfs are $f_{W_{i}}(w)=\frac{w^{\nu_{i}/2-1}\,e^{-w/2}}{2^{\nu_{i}/2}\Gamma(\nu_{i}/2)}$, $w>0$, $i=1,2$. There is no closed-form pdf for the weighted sum $\sqrt{gW_{1}/\nu_{1}+hW_{2}/\nu_{2}}$; this is the core obstacle of the Behrens–Fisher problem, so we keep the bivariate $(W_{1},W_{2})$ pair intact.
@@ 2.1 MB expansion of the square-root factor
$(x+y)^{\frac{1}{2}}=\frac{1}{2\pi i}\int\limits_{c-i\infty}^{c+i\infty}\frac{\Gamma(-s)\,\Gamma(\tfrac{1}{2}+s)}{\Gamma(\tfrac{1}{2})}x^{s}\,y^{1/2-s}\,ds.$
@@
The inseparable factor is now a weighted geometric mean $(g/\nu_{1})^{s}w_{1}^{s}(h/\nu_{2})^{1/2-s}w_{2}^{1/2-s}$ multiplied by a pures-s coefficient (the $\Gamma$-product). All $w_{1}$-dependence and $w_{2}$-dependence are split into separate powers. This makes the forthcoming integrations elementary. Since absolutely convergent integrals may be rearranged, we move the MB contour outside:
@@
where $\alpha=\dfrac{t^{2}}{2(g+h)}$. With this development, (7) is precisely the single-contour formula. Two dimensions have become one; the price is the integral over a vertical line in the complex plane. From here, residue calculus yields an explicit series by closing the contour to the right: poles of $\Gamma(-s)$ at $s=0,1,\dots$ give the hypergeometric-type series, which terminates when either $\nu_{1}$ or $\nu_{2}$ is odd. Alternatively, before the MB step, one may collapse $(w_{1},w_{2})$ to $(S,U)$. But, the MB path is shorter: from (7), shift the contour left, pick the poles of $\Gamma(\tfrac{1}{2}+s)$, and recognize the resulting Beta function integral as Gauss’ hypergeometric, thereby facilitating a compact expression for $f_{T}(t)$ that becomes apparent in the sequel. More importantly, the integrand in (7) is exactly of the $\gamma$-product type that RMT turns into asymptotic series without further work: expand the $\Gamma$-ratio around large $s$ to read off the $\lvert t\rvert\to\infty$ expansion coefficients directly, thereby yielding exact tail coefficients.
The MB approach works for all $g$, $h$. While Taylor series in $w_{1}/w_{2}$ converges only when one weight is smaller than the other, the MB integral is valid everywhere. It handles non–integer exponents naturally. In our case, the exponent is $\rho=\tfrac{1}{2}$; the MB representation is suitable for arbitrary complex $\rho$. Note that, MB aids interfacing with $\Gamma$-distributions: $\chi^{2}$ densities are special cases of $\Gamma$-densities. The MB splits $(x+y)^{\rho}$ into $\Gamma$-functions of $s$ times monomials in $x$ and yy—perfectly aligned with $\Gamma$ integrals. Furthermore, the MB integrals provide contour flexibility. We may choose to close left or right, whichever set of poles yields the nicer series (terminating, fastest convergence, etc).
@@ 2.2 Residue calculus on the MB contour
$\Gamma(-s)$ $s=0,1,2,\dots$ Right-hand ladder.
$\Gamma(\tfrac{1}{2}+s)$ $s=-\tfrac{1}{2},-\tfrac{3}{2},-\tfrac{5}{2},\dots$ Left-hand ladder.
$\Gamma\left(s+\tfrac{\nu_{1}}{2}\right)$ $s=-\tfrac{\nu_{1}}{2},-\tfrac{\nu_{1}}{2}-1,\dots$ Left ladder, terminates if $\nu_{1}$ even.
$\Gamma\left(\tfrac{\nu_{2}}{2}-s+\tfrac{1}{2}\right)$ $s=\tfrac{\nu_{2}}{2}+\tfrac{1}{2}+k,\ k=0,1,\dots$ Right ladder, terminates if $\nu_{2}$ odd.
@@
the factor $\Gamma\!\left(\tfrac{\nu_{2}}{2}-k+\tfrac{1}{2}\right)$ vanishes once $k>\tfrac{\nu_{2}}{2}-\tfrac{1}{2}$ if $\nu_{2}$ is an odd integer because the $\Gamma$-argument hits a non-positive integer. The same happens to $\Gamma\!\left(k+\tfrac{\nu_{1}}{2}\right)$ when $\nu_{1}$ is even and we close the contour to the left (alternative route). Thus the series terminates whenever either degrees of freedom parameter is odd—a property first noted by Fisher (1935). Collecting all right-hand residues and the front constants in (10) gives
@@
This series is valuable for the following reasons: The $\Gamma$-coefficients are rational in $\nu_{1},\nu_{2}$; algebra systems can integrate or differentiate $f_{T}(t)$ term-wise to obtain moments, mgf, cdf expansions, etc. Only a handful of terms are needed for double-precision accuracy when $t\lesssim 10$. Truncation after $m$ terms leaves a remainder $O(t^{-(\nu_{1}+\nu_{2}+1)+2m})$, giving controlled error bounds directly. Recognizing the ratio of $\Gamma$’s as Beta coefficients, (12) rearranges into a Gauss hypergeometric series $_{2}F_{1}$; summation leads to a closed form. If $g=h$, the factor $(g/h)^{k}$ becomes 1 and the $\Gamma$-triple collapses to a single $\Gamma$-ratio, giving the $t_{\nu_{1}+\nu_{2}}$ density.
Summing (12) with the Euler-type identity for $_{2}F_{1}$ yields compact expression for $f_{T}(t)$ which we will provide in the next section. That closed form is preferred for large $\lvert t\rvert$, because the hypergeometric behaves smoothly, while the plain series may suffer cancellation. The same $\Gamma$-product that gives each residue also feeds Ramanujan’s master theorem. Essentially, we write $\Psi(s)\,\Gamma(s+\nu_{1}/2)\,\Gamma(\nu_{2}/2-s+1/2)\equiv\Gamma(s)\,A(-s)$, and read off the coefficients $A(m)$ directly for the $|t|\to\infty$ expansion. When $\nu_{i}$, $i=1,2$, are both odd, we get a finite series that leads to an elementary rational function using duplication. When $\nu_{i}$ are both even, we get a hypergeometric series of type $_{2}F_{1}$ with integer parameters, reducible further to complete elliptic integrals for certain balanced sample-size ratios. Reducing the contour integral into residues is central to our work that transforms an abstract complex integral into an explicit computational tool. It reveals hidden algebraic cancellations, bridges to standard special-function framework (hypergeometric/Beta) and prepares the ground for both numerical evaluation and asymptotic analysis. Everything that follows—closed-form hypergeometric, tail coefficients, moment formulas—starts with the series (12). Without identifying and summing the $\Gamma(-s)$ residues, the Behrens–Fisher density would remain locked inside the contour integral; after residual analyses, it becomes fully transparent.
@@
The MB approach in the previous sections already gave a series. However, series are not always the best numerical object—convergence slows in the tails and severe cancellation can occur when $g\approx h$. A single real integral with Euler–Beta weight can sometimes be integrated analytically, producing a compact special-function formula (see (Andrews et al., 1999, Section 2.3)). Most symbolic integrators (Mathematica, Maple, etc.) recognize Euler-type kernels and instantly reduce them to the classical Gauss function $_{2}F_{1}$. Hence, we reorganize the starting double integral so that all dependence on one variable disappears, leaving a 1-D integral of Euler–Beta form.
@@
where $\mathcal{C}$ is the set of all constants independent of $U$: $\mathcal{C}=\frac{\Gamma(c)(g/\nu_{1})^{-a}(h/\nu_{2})^{-b}}{2^{(a+b)}\Gamma(a)\Gamma(b)\sqrt{2\pi(g+h)}}$, with $a=\frac{\nu_{1}}{2}$, $b=\frac{\nu_{2}}{2}$, $c=\frac{\nu_{1}+\nu_{2}+1}{2}$. Upon closer inspection, we can write, with $\lambda=\frac{\nu_{1}h}{2\nu_{2}g}-\frac{\alpha h}{\nu_{2}}-1$, $f_{T}(t)=\mathcal{C}(\nu_{2}/h)^{-c}\int_{0}^{1}\!u^{a-1}(1-u)^{b-1}(1+\lambda u)^{-c}du$.
By Euler’s integral representation of Gauss’ hypergeometric function (see (Abramowitz and Stegun, 1964, Chapter 15)), for $\Re c>0$, $\int_{0}^{1}\!u^{a-1}(1-u)^{b-1}(1+\lambda u)^{-c}du=B(a,b){}_{2}F_{1}(c,a;a+b;-\lambda)$, with $B(a,b)=\Gamma(a)\Gamma(b)/\Gamma(a+b)$ so that the Beta factor cancels the $\Gamma(a)\Gamma(b)$ already in $\mathcal{C}$. Letting $\rho(t)=\frac{g^{-1}-h^{-1}}{h^{-1}+t^{2}/(g+h)}$, and organizing the terms together yields
@@
where $\rho(t)$ is the $t$-adjusted version of $\lambda$ that appears in the fully scaled hypergeometric formula.
@@
Natural starting point for Ramanujan’s theorem/saddle-point asymptotics Immediate access to Gauss–Kummer transformations (yielding the hypergeometric functions) for numerical conditioning.
@@
It is worth noting that the MB representation and Euler-type integral collapses represent two analytically distinct yet ultimately equivalent approaches to deriving the same result. For symbolic series manipulations, MB is the preferred approach. For compact numeric evaluation, Euler–hypergeometric is preferable. Table 2 summarizes the differences between these two approaches. Scientific computing libraries (e.g., SciPy) support high-precision evaluation of $_{2}F_{1}$ with rigorous error control, enabling accurate computation of $f_{T}(t)$ for any parameters $(\nu_{1},\nu_{2},g,h,t)$. Interchanging the two samples, i.e., $(g,\nu_{1})\leftrightarrow(h,\nu_{2})$, leaves $f_{T}(t)$ invariant. This follows from the symmetry of $_{2}F_{1}(c,a;a+b;-\rho)$ in $(a,b)$ when $\rho\to-\rho$. In the equal-variance limit, as $g\to h$, we have $\rho\to 0$ and $_{2}F_{1}\to 1$, so (16) reduces to the Student $t$ density with $\nu_{1}+\nu_{2}$ degrees of freedom. In the large–sample limit, as $\nu_{1},\nu_{2}\to\infty$, the distribution converges to $\mathcal{N}(0,1)$. The asymptotic behavior of $_{2}F_{1}(c,a;a+b;-z)$ as $z\to\infty$ follows directly from Euler’s integral representation, yielding an $O\left(t^{-(\nu_{1}+\nu_{2}+1)}\right)$ decay. This matches the tail derived via complex–residue analysis. Term–by–term integration of (16) in $t$, under the substitution $t=\sqrt{g+h}\tan\theta$, converts the expression into a sum of Appell $F_{1}$ functions—amenable for numeric evaluation. The one-dimensional collapse is the key that unlocks a single Gauss hypergeometric function for $f_{T}$. This form encapsulates all parameter dependence in five $\Gamma$’s and one $_{2}F_{1}$, is suitable for standard libraries and asymptotic expansions, and shows that the Behrens–Fisher density lives in the classical hypergeometric family, explaining why equal-variance data drop to Student-$t$ (the $_{2}F_{1}\!\equiv 1$ case) and why odd degrees of freedom collapse to elementary expressions (because the Gauss function terminates). Subsequent steps—cdf calculation, saddle-point corrections, or benchmarking Welch–Satterthwaite—benefits from having this compact representation.